# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@ 4.1. Reduction theorem for Seifert manifolds
(4.5) $\sum_{h\in H}f_{h\sigma^{\text{ac}}}(t)h=t^{(k-2)A-\sum_{i}A_{i}}\frac{(g_{1}t^{2A_{1}}-1)\cdots(g_{k}t^{2A_{k}}-1)}{(g_{0}t^{2A}-1)^{k-2}}.$
Proof.
We will start with the series $Z_{0}(q)$ for simplicity. The idea is to reduce the multivariate integrand of $Z_{0}$ to univariate rational function $f_{0}(t)$, and then study the effect of this operation on the exponents given by the theta function. The function $f_{0}(t)$ is obtained from the integrand in $Z_{0}(q)$ by the following substitution:
(4.6) $\begin{array}[]{ll}z_{i}\,\rightarrow t^{A_{i}}&\mbox{for the leaves}\\ z_{0}\rightarrow t^{A}&\mbox{for the node.}\\ \end{array}$
When computing $Z_{0}(q^{|H|})$ from the definition, we substitute each monomial $\vec{z}\,^{\vec{l}}$ with $q^{(\vec{l},\operatorname{adj}(-M)\vec{l})/4}.$ We can expand the exponent (ignoring factors of 4 in what follows):
(4.7) $(\vec{l},\operatorname{adj}(-M)\vec{l})=l_{0}^{2}A+2\sum_{i\neq 0}l_{0}l_{i}A_{i}+\sum_{i\neq j}l_{i}l_{j}A_{ij}+\sum_{i\neq 0}l_{i}^{2}\operatorname{adj}(-M)_{ii}.$
(4.8) $(l_{0}A+\sum_{i\neq 0}l_{i}A_{i})^{2}/A$ $=l_{0}^{2}A+2\sum_{i\neq 0}l_{0}l_{i}A_{i}+\sum_{i,j\neq 0}l_{i}l_{j}A_{i}A_{j}/A$
(4.9) $=l_{0}^{2}A+2\sum_{i\neq 0}l_{0}l_{i}A_{i}+\sum_{i,j\neq 0}l_{i}l_{j}A_{ij}.$
We see that the two expressions (4.7) and (4.9) differ only in terms corresponding to “a leaf with itself”, $l_{i}^{2}A_{ii}$ vs. $l_{i}^{2}\operatorname{adj}(-M)_{ii}$. Since $l_{i}^{2}=(\pm 1)^{2}=1$ for each vector $\vec{l}\in\mathbb{Z}^{s}$ that contributes a nonzero term to $Z_{0}(q)$, these numbers give the same contribution to every term of the sum, so they contribute by an overall power of $q$ in the series $Z_{0}(q).$
The overall $q$-power correction $\Lambda$ is
$=-|H|(3s+\text{Tr}(M)+\sum_{i}(2-\delta_{i})(M^{-1})_{ii}-(2-\delta_{0})(M^{-1})_{00})-\frac{1}{A}\sum_{i\neq 0}A_{i}^{2}$
$=24\lambda(Y)-A(2-\delta_{0}+\sum_{i\neq 0}\frac{1}{a_{i}^{2}})$
using the formula Eq. 3.7 for Casson–Walker invariant $\lambda$ as in the proof of Theorem 3.5. To obtain the formula for $\widehat{Z}_{\sigma}(q)$, we follow the same computation of the exponents (without scaling $q$ by $|H|$), applied on the equivariant expression (2.7). ∎
The reduction theorem can be considered an analog of the reduction procedure of Némethi [47, p. 364], which was developed in greater generality. It is natural to proceed in this direction for more general graphs and try to obtain more general reduction theorems for $\widehat{Z}_{\sigma}(q)$ invariants. A clue toward a more general reduction was already presented in [6] for $H$-shaped graphs with exactly 6 nodes. Splice diagrams provide a natural framework for generalizing these computations for any plumbing.
A useful corollary of Theorem 4.2 explains frequent vanishing of the $\widehat{Z}_{\sigma}(q)$ series for Seifert manifolds:
Corollary 4.3.
Let $Y$ be a Seifert manifold with $k$ singular fibers as above, for which the element $g_{0}\in H_{1}(Y,\mathbb{Z})$ is trivial, i.e. $|e|\text{lcm}(a_{1},\dots,a_{k})=1$. Then there are at most $2^{k}$ nonzero $\widehat{Z}_{\sigma}$ invariants.
Proof.
As $g_{0}$ vanishes, the only elements of $H$ which appear in the expansion of
(4.10) $\sum_{h}f_{h\cdot\sigma^{\text{ac}}}(t)=t^{(\dots)}\frac{(g_{1}t^{2A_{1}}-1)\cdots(g_{k}t^{2A_{k}}-1)}{(g_{0}t^{2A}-1)^{k-2}}$
are of the form
$h=g_{1}^{u_{1}}g_{2}^{u_{2}}\cdots g_{k}^{u_{k}},$
where $u_{i}\in\{0,1\}$, so there is $2^{k}$ of them. The nonzero $f_{h\cdot\sigma^{\text{ac}}}$ are their coefficients. ∎
@@
For any positive integer $n$, Seifert manifold $M(2;(n,n-1),(n,n-1),(n,1))$ has $|H|=n^{3}/n=n^{2}$ and $|g_{0}|=n/n=1$ so there are only eight nonzero $q$-series $\widehat{Z}_{\sigma}(q).$ On the other hand, consider $M(1;(3,1),(4,1),(5,1))$ [^11] The link of $S_{12}$ singularity $x^{3}y+y^{2}z+xz^{2}=0.$. It has $|H|=13$ and all thirteen $\widehat{Z}_{\sigma}(q)$ are nonzero, as $g_{0}$ generates $H.$
Example 4.5.
Let us illustrate in detail how the reduction theorem works on the Seifert manifold $Y=M(2;(2,1),(3,2),(3,2))$, the link of $E_{6}$ singularity $x^{2}+y^{3}+z^{4}=0$. A plumbing graph and splice diagram for $Y$ are given in Fig. 8.
Figure 8. Plumbing graph and splice diagram of Seifert manifold $Y=M(2;(2,1),(3,2),(3,2))$
$Z_{0}(q)$: Consider the function $f_{0}$:
(4.11) $f_{0}(t)=\frac{(t^{9}-t^{-9})(t^{6}-t^{-6})(t^{6}-t^{-6})}{t^{18}-t^{-18}}=\frac{(t^{9}-t^{-9})(t^{12}+t^{-12}-2)}{t^{18}-t^{-18}}.$
$\frac{1}{2}(\dots+t^{-33}+2t^{-27}+t^{-21}-t^{-15}-2t^{-9}+t^{-3}+t^{3}-2t^{9}-t^{15}+t^{21}+2t^{27}+t^{33}+\dots)$
Laplace transform $\mathcal{L}_{18}$ substitutes each monomial $t^{n}$ with $q^{n^{2}/{4\cdot 18}}$:
$q^{1/8}(1-2q-q^{3}+q^{6}+2q^{10}+q^{15}-q^{21}-2q^{28}-q^{36}+q^{45}+2q^{55}\dots=q^{25/8}Z_{0}(q^{3}).$
The exponent $25/8$ is related to the Casson–Walker invariant $\lambda(Y)=-11/12$ as in the proof:
$-\frac{25}{8}=-\Lambda=6\lambda(Y)+\frac{18}{4}-\frac{6^{2}+6^{2}+9^{2}}{4\cdot 18}.$
$\widehat{Z}_{\sigma}(q)$: We have $e=-1/6$ and $|H|=|e|a_{1}a_{2}a_{3}=3$. The group $H$ is generated by $g_{1},g_{2},g_{3},g_{0}$ satisfying $g_{1}^{2}=g_{0}$, $g_{2}^{3}=g_{0}$, $g_{3}^{3}=g_{0},g_{1}g_{2}^{2}g_{3}^{2}=g_{0}^{2}$. It is easy to see that $g_{0}=g_{1}=1$, and put $g:=g_{2}=g_{3}^{2}$ and $\sigma:=\sigma^{\text{ac}}$. The rational functions in Theorem 4.2 read
$\sum_{h}f_{h\cdot\sigma}(t)h$ $=t^{18-9-6-6}\frac{(g_{1}t^{18}-1)(g_{2}t^{12}-1)(g_{3}t^{12}-1)}{g_{0}t^{36}-1}$
$=t^{-3}\frac{(gt^{12}-1)(g^{2}t^{12}-1)}{1+t^{18}}$
$=\frac{t^{21}+t^{-3}}{1+t^{18}}+(g+g^{2})\frac{-t^{9}}{1+t^{18}}$
These functions are expanded as
$\operatorname{s.e.}f_{\sigma}(t)$ $=\frac{1}{2}(\dots+t^{-21}-t^{-15}+t^{-3}+t^{3}-t^{-15}+t^{21}\dots$
$\operatorname{s.e.}f_{g\sigma}(t)=\operatorname{s.e.}f_{g^{2}\sigma}$ $=\frac{1}{2}(\dots-t^{-45}+t^{-27}-t^{-9}-t^{9}+t^{27}-t^{45}\dots$
After performing the Laplace transform $\mathcal{L}_{A|H|}:t^{n}\to q^{n^{2}/4\cdot 18\cdot 3}$ and multiplying by $q^{\Lambda/3}=q^{25/24}$ we obtain the $q$-series:
$\widehat{Z}_{\sigma}(q)$ $=q^{-1}-1+q+q^{4}-q^{6}-q^{11}+q^{14}+q^{21}-q^{25}-q^{34}+q^{39}+q^{50}-q^{56}-\dots$
$\widehat{Z}_{g\sigma}(q)$ $=\widehat{Z}_{g^{2}\sigma}(q)=q^{-2/3}(-1+q^{3}-q^{9}+q^{18}-q^{30}+q^{45}-q^{63}+q^{84}-q^{108}+\dots.$
4.2. Lens spaces
As another application of Theorem 4.2, we compute $\widehat{Z}_{\sigma}(q)$ of lens space $Y=L(p,r)=M(0;(r,p),(1,0),(1,0))$, which has been done systematically only in the case of $r=1$ [33]. We get:
$Z_{0}(Y;q^{p})$ $=q^{\Lambda}\mathcal{L}_{r}\left(\frac{(t-t^{-1})(t^{r}-t^{-r})^{2}}{(t^{r}-t^{-r})}\right)$
$=2q^{6\lambda(Y)}q^{-r/4-1/4r}\mathcal{L}_{r}(t^{r+1}-t^{r-1})$
$=2q^{6\lambda(Y)}(q^{1/2}-q^{-1/2})$
$=q^{6\lambda(Y)}Z_{0}(S^{3},q)$
We see that $Z_{0}(Y;q^{p})$ is proportional to a $q$-power [^12] Casson–Walker invariant $\lambda(Y)$ equals $p\cdot s(r,p)/2$ where $s(r,p)$ is the Dedekind sum [47, Ex. 9.3.6]. of $\lambda(Y)$, with a fixed factor being exactly $Z_{0}(q)$ of the 3-sphere $S^{3}$. As $S^{3}$ is the universal abelian cover of the lens space, we get a direct relation between the invariants of the base space and the cover.
For the equivariant version, we have
(4.12) $\sum_{h\in H}\widehat{Z}_{h\sigma^{\text{ac}}}(q)h$ $=q^{\Lambda}\mathcal{L}_{b}(t^{-1-r}(g_{1}t^{2}-1)(g_{0}t^{2r}-1))$
(4.13) $=q^{6\lambda(Y)/p}((g_{0}g_{1}+1)q^{1/2p}-(g_{0}+g_{1})q^{-1/2p})$
(4.14) $=q^{6\lambda(Y)/p}((g^{r+1}+1)q^{1/2p}-(g^{r}+g)q^{-1/2p})$
where $g=g_{1}$ and $g_{0}$ are the generators of $H$ satisfying $g_{1}^{r}=g_{0}$ and $g_{1}^{p}=1.$
An interesting corollary of this computation is that $\widehat{Z}(q)$ invariants distinguish lens spaces. Indeed, the expression 4.12 is a version of $q$-deformed Reidemeister–Turaev torsion (as conjectured in [12]), which is known to have this property. It is interesting to notice that one needs to consider the whole package of $\widehat{Z}_{\sigma}(q)$ invariants — the set of exponents of $q$ itself does not distinguish lens spaces.
5. $Z_{0}$, irreducible connections, spectrum
In this last section, we explore an interplay between four different objects associated with a Seifert manifold $Y$: a generalized version of $Z_{0}$ $q$-series, irreducible flat $\mathop{SL}(2,\mathbb{C})$-connections, the spectrum of the associated splice type singularity, and the related vertex algebras. In Section 3 we examined which invariants of the 3-manifold depend on the full data of the plumbing graph, or rather just on the splice diagram. With the exception of vertex algebras, all objects listed above are of the latter type. [^13] In particular, all of them become essentially trivial for lens spaces. It is an interesting question whether the discussion in this section generalizes to more complicated plumbed manifolds.
5.1. Irreducible flat connections on Seifert manifolds
The study of flat connections on 3-manifolds (equivalently, representations of the fundamental group), is an important tool in 3-dimensional topology. Flat connections are critical points of the Chern-Simons functional, so they are expected to organize the structure of both perturbative and non-perturbative complex Chern-Simons theory [29]. With $\widehat{Z}_{\sigma}(q)$ providing a non-perturbative definition of this theory, the role of complex flat connections on $Y$ extended further to the category of line operations, dubbed $\text{MTC}[Y,G_{\mathbb{C}}]$ in [35].
We continue to work with the simplest non-trivial choice of the complex gauge group, $G_{\mathbb{C}}=\mathop{SL}(2,\mathbb{C})$, and leave generalizations to an interested reader and future work. Let us denote by $\mathcal{M}^{*}(Y,\mathop{SL}(2,\mathbb{C}))$ the moduli space of irreducible flat connections on $Y$ (modulo conjugation). For Seifert homology spheres, the moduli space can be described as a certain moduli space of parabolic Higgs bundles over the (orbifold) base $\mathbb{C}\mathbb{P}^{1}$ [5]. The moduli space is smooth, noncompact, and has components of various complex dimensions between 0 and $62k-$ 6, where $k$ is the number of singular fibers. In particular, for three singular fibers, it is just a collection of points. A component $C\subset\mathcal{M}^{*}(Y,\mathop{SL}(2,\mathbb{C}))$ of dimension $62n-$ 6 has Euler characteristics
(5.1) $\chi(C)=(n-1)(n-2)2^{n-4}=\binom{n-1}{2}2^{n-3},$
which is also equal to the rank of the cohomology of $C$ as it is nontrivial only in even degrees.
Related to this, $\mathop{SL}(2,\mathbb{C})$-Casson invariant $\lambda^{C}_{\mathop{SL}(2,\mathbb{C})}(Y)$ was defined in [16]. For Seifert homology spheres, it simply counts the zero-dimensional components [4]. An alternative definition, denoted by $\lambda^{P}_{SL(2,\mathbb{C})}(Y)$, is due to Abouzaid and Manolescu [1], which in the case of Seifert homology spheres coincides with the Euler characteristics of the moduli space of flat connections. The number of components of dimension $62n-$ 6 is
(5.2) $\frac{e_{n}(a_{1}-1,a_{2}-1,\dots,a_{k}-1)}{2^{n-1}}$
where $e_{n}$ is the elementary symmetric polynomial
$e_{n}(x_{1},x_{2},\dots,x_{k})=\sum_{1\leq j_{1}<j_{2}<\dots<j_{k}\leq n}x_{j_{1}}x_{j_{2}}\dots x_{j_{k}}.$
For example, for $Y=\Sigma(2,3,5,7)$ we have $e_{3}(1,2,4,6)/4=23$ zero-dimensional components and $e_{4}(1,2,4,6)/8=6$ two-dimensional components and
$\lambda^{P}_{SL(2,\mathbb{C})}(Y)=23+6\cdot 6=59.$
The set of components of dimension $62n-$ 6 can be identified with the set $D$ of tuples of integers $(d_{1},d_{2},\dots,d_{k})$ satisfying $0<d_{i}<a_{i}$ and exactly $n$ numbers being nonzero, modulo the action of $(\mathbb{Z}_{2})^{n-1}$ that takes $d_{i}\rightarrow a_{i}-d_{i}$ simultaneously for an even number of nonzero $d_{i}$, resulting in the count above. The numbers $d_{i}$ can be thought of as rotational numbers of the corresponding representation with the choice of a representant in the $(\mathbb{Z}_{2})^{n-1}$ orbit corresponding to taking different presentations of the group. For more details, see [64] for the case of 3 singular fibers and [3] for any number of fibres. [^14] Note that some components of lower dimensions were omitted there due to a too restrictive fundamental domain for the $(\mathbb{Z}_{2})^{n-1}$ action.
5.2. From connections to spectrum
Recall from Section 3, that to some classes of plumbed 3-manifolds, one can naturally assign (an equisingular family of) splice-type singularities defined using the splice diagram. This is particularly nice for Seifert manifolds, where the corresponding singularity is a complete intersection of Brieskorn type [51].
We will start with three singular fibers for simplicity. Given a Seifert manifold $Y=M((a_{1},b_{1}),(a_{2},b_{2}),(a_{3},b_{3}))$, the corresponding splice-type singularity $X$ is
(5.3) $x^{a_{1}}+y^{a_{2}}+z^{a_{3}}=0.$
By Neumann’s theorem, this is the universal abelian cover of the quasihomogeneous singularity with link $Y$. The equation is constructed from the splice diagram and does not depend on $b_{i}$.
This singularity, being a hypersurface, has natural one-parameter smoothing and admits the monodromy operator and spectral numbers defined using a canonical mixed Hodge structure on the Milnor fiber. The rank of the cohomology of the Milnor number (the number of spectral numbers with multiplicities) is Milnor number $\mu.$
The spectrum of the singularity (5.3) is computed using Thom–Sebastiani theorem. It is simply the set
(5.4) $\frac{d_{1}}{a_{1}}+\frac{d_{2}}{a_{2}}+\frac{d_{3}}{a_{3}},\quad 0<d_{i}<a_{i}\text{ for }i=1,2,3$
The eigenvalues of the monodromy are equal to exponentiated spectral numbers $\exp(2\pi i\sum\frac{d_{i}}{a_{i}})$. From (5.4) and (5.2) we see that, in the case of $Y$ being homology sphere, the Milnor number of $X=X^{ab}$ is proportional to $\lambda^{C}_{SL(2,\mathbb{C})}(Y)=\lambda^{P}_{SL(2,\mathbb{C})}(Y)$; see [4]:
(5.5) $\frac{\mu(X^{ab})}{4}=\lambda^{C}_{SL(2,\mathbb{C})}(Y),$
If $Y$ is a rational homology sphere, the left-hand side of (5.5) needs to be replaced by the number of $(\mathbb{Z}/2\mathbb{Z})^{2}$ orbits (the action now has fixed points). The count of connection was given in [15, p. 20] and the analog of the equation above follows easily from Burnside’s lemma.
This relation extends nicely to more singular fibers, as we will describe now. Although this extension may be known to the experts, we could not find it in the literature. We will formulate it in terms of the monodromy eigenvalues as those were explicitly computed by Hamm [37]. The spectrum of complete intersections has been defined [18], but we are now merely interested in the counts of the elements, so it does not make a big difference.
Let $Y=M((a_{1},b_{1}),\dots,(a_{k},b_{k}))$ be a Seifert manifold with $k$ singular fibres. The corresponding splice-type equations are a complete intersection of $k-2$ equations of type (5.3) in $k$ variables, again with generic coefficients. More specifically, let us take a complex matrix $\alpha_{ij}$ of size $(k-2)\times k$ with all maximal subdeterminants being nonzero. Consider the functions $f_{i}(x):\mathbb{C}^{k}\rightarrow\mathbb{C},i=1\dots k-2$ given by
(5.6) $f_{i}(x)=\sum\alpha_{ij}x_{j}^{a_{j}}$
The set
$f_{1}(x)=f_{2}(x)=\dots=f_{k-2}(x)=0$
defines a complex surface $X$ with an isolated complete intersection singularity at the origin. If $Y$ is a homology sphere, then $Y$ is the link of $X$ [51]. If it is a rational homology sphere, the link of $X$ is $Y^{ab}$.
We will now assume that $Y$ is a homology sphere for simplicity. [^15] This is the case mostly studied in the literature. We expect that for rational homology spheres, similar modifications can be done as in the case of 3 singular fibers. If we now take as $X^{*}$ the surface given by the equations (5.6) but the last one $f_{k-2}$, the function $f_{k-2}$ is regular on $X^{*}$ except 00, and we can define the monodromy operator $h$ of the Milnor fiber of $f_{k-2}$. We have
Lemma 5.1 (Hamm).
The characteristic polynomial $\chi(t)$ of $h$ is given by
(5.7) $\chi(t)=\prod_{\begin{subarray}{c}3\leq n\leq k,\\ 1\leq j_{1}<\dots<j_{n}\leq k\end{subarray}}\delta_{a_{j_{1}},a_{j_{2}},\dots,a_{j_{n}}}(t)^{\binom{n-1}{2}}$
where
(5.8) $\delta_{c_{1},c_{2},\dots,c_{n}}(t)=\prod_{\begin{subarray}{c}d_{1},\dots,d_{n},\\ 1\leq d_{i}<c_{i},\\ i=1,\dots,n\end{subarray}}e^{2\pi i(d_{1}/c_{1}+d_{2}/c_{2}+\cdots+d_{n}/c_{n})}-t$
We see that the eigenvalues of monodromy are labeled by tuples $(d_{1},d_{2},\dots,d_{n})$ for $n$ varying from 3 to $k$ in the same way as the $62n-$ 6 dimensional components of the moduli space of connections. The multiplicity of the eigenvalue is $\binom{n-1}{2}$ so for each $n$, we get that the number of eigenvalues (with multiplicities) is four times the number of $62n-$ 6 dimensional components times the Euler characteristics $\binom{n-1}{2}2^{n-3}$. We obtain
Proposition 5.2.
Let $Y$ be a Seifert homology sphere and $X$ the corresponding complete intersection singularity with Milnor number $\mu(X)$. Then
$\lambda^{P}_{SL(2,\mathbb{C})}(Y)=\frac{\mu(X)}{4}.$
The above considerations suggest that this relation is not just a numerical coincidence and should be given an interpretation on the level of cohomology. Abouzaid and Manolescu defined the cohomology using Heegaard splitting, giving two Lagrangians in the character variety of the Heegaard surface. They considered certain perverse sheaf of vanishing cycles, which can be thought of as being associated with a function related to the Lagrangians. [^16] In geometric and physics realizations of knot and 3-manifold invariants this function is known as the superpotential of a 3d theory [56, 25, 26, 32, 21, 20, 62]; it defines a Landau-Ginzburd (LG) model that we discuss in more detail shortly. The resulting cohomology is expected to be closely related to $\text{MTC}[Y,G_{\mathbb{C}}]$ and to the Floer homology of Vafa-Witten theory on $Y$; see e.g. [27, 35, 36]. In particular, these connections indicate a direct relation to the sheaf counting (Vafa-Witten theory) on the complex surface $X$.
Further support for this relation comes from the analogy between the above proposition and the “Casson invariant conjecture” relating the $\lambda_{SU(2)}$ with the one eight of the signature $\sigma$ of the Milnor fiber [53]. Since the latter has been verified for splice-quotients in [49], one might expect a suitable $\mathop{SL}(2,\mathbb{C})$ analogue to hold true in the same generality as well.
5.3. Nonabelian noncommutative Hodge Structure
Consider a germ of a singular affine hypersurface $X$ (more generally, a complete intersection) with an isolated singular point 00, with the link $Y$ of $X$ being a homology sphere. [^17] For rational homology spheres, one would have to pass to the universal abelian covering $X^{ab}$, which makes the discussion more involved. We have
(5.9) $\pi_{1}(X{\setminus}0)\cong\pi_{1}(Y)$
so we can use the algebraic structure on $X{\setminus}0$ to study moduli spaces of representations.
Following C. Simpson [66] and T. Mochizuki we have (see the survey of Sabbah [63]), we have a real-analytic isomorphism of the Betti and Dolgachev moduli spaces.
(5.10) $H^{1}_{B}(Y)\cong M_{Dol}(X{\setminus}0).$

### v3

@@ 5.1.2 Seifert homology spheres
$\delta_{c_{1},c_{2},\dots,c_{n}}(t)=\prod_{\begin{subarray}{c}d_{1},\dots,d_{n},\\ 1\leq d_{i}<c_{i},\\ i=1,\dots,n\end{subarray}}\bigl{(}{\rm e}^{2\pi{\rm i}(d_{1}/c_{1}+d_{2}/c_{2}+\cdots+d_{n}/c_{n})}-t\bigr{)}.$
We see that the eigenvalues of $h$ are labeled by tuples $(d_{1},d_{2},\dots,d_{n})$ for $n$ from 3 to $k$ in the same way as are the (2n6)(2n-6)-dimensional components of $\mathcal{M}^{*}(Y,\operatorname{SL}_{2}(\mathbb{C}))$ (without the $(\mathbb{Z}_{2})^{n-1}$ action). Moreover, the eigenvalues have multiplicities $\binom{n-1}{2}$. Together, the Milnor number is
$\mu(X)=\sum_{3\leq n\leq k}e_{n}(a_{1}-1,\dots,a_{k}-1)\binom{n-1}{2}$,
which gives the result by comparing with (5.10). ∎
@@
For $Y=\Sigma(2,3,5,7)$, we have $e_{3}(1,2,4,6)/4=23$ zero-dimensional components and $e_{4}(1,2,4,6)/8=6$ two-dimensional components and $\lambda^{P}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)=23+6\cdot 6=59$. The corresponding Brieskorn–Hamm singularity $X$ has 92 eigenvalues of multiplicity 1 and 48 eigenvalues of multiplicity 3, so that $\mu(X)=236=4\cdot 59$.
The above considerations suggest that this relation is not just a numerical coincidence and should be given an interpretation on the level of cohomology of the moduli space $\mathcal{M}^{*}(Y,\operatorname{SL}_{2}(\mathbb{C}))$ and of the Milnor fiber. Abouzaid and Manolescu defined the sheaf-theoretic $\operatorname{SL}_{2}(\mathbb{C})$ Floer homology using Heegaard splitting, giving two Lagrangians in the character variety of the Heegaard surface. They considered certain perverse sheaf of vanishing cycles, which can be thought of as being associated with a function related to the Lagrangians. [^15] In geometric and physics realizations of knot and 3-manifold invariants this function is known as the superpotential of a 3d theory [20, 21, 25, 26, 31, 56, 62]; it defines a Landau–Ginzburg (LG) model that we discuss in more detail shortly. The resulting cohomology is expected to be closely related to $\text{MTC}[Y,G_{\mathbb{C}}]$ and to the Floer homology of Vafa–Witten theory on $Y$; see, e.g., [27, 34, 35]. In particular, these connections indicate a direct relation to the sheaf counting (Vafa–Witten theory) on the complex surface $X$.
Further support for this relation comes from the analogy between the above proposition and the “Casson invariant conjecture” relating the $\lambda_{\operatorname{SU}(2)}$ with the one eight of the signature $\sigma$ of the Milnor fiber [53]. Since the latter has been verified for splice-quotients in [49], it would be interesting to explore a suitable analogue for $\operatorname{SL}_{2}(\mathbb{C})$.
5.2 Nonabelian noncommutative Hodge structure (NCNA)
In this section, we give some questions and speculative interpretations of the relation between the moduli spaces of flat connections and the spectra of singularities. We will develop this further in separate paper.
Consider a germ of a singular affine hypersurface $X$ (more generally, a complete intersection) with an isolated singular point 0, with the link $Y$ of $X$ being a homology sphere. [^16] For rational homology spheres, one would have to pass to the universal abelian covering $X^{\rm ab}$, which makes the discussion more involved. We have $\pi_{1}(X{\setminus}0)\cong\pi_{1}(Y)$, so we can use the algebraic structure on $X{\setminus}0$ to study moduli spaces of representations.
Following C. Simpson [65] and T. Mochizuki (see the survey of Sabbah [63]), we have a real-analytic isomorphism of the Betti and Dolgachev moduli spaces $H^{1}_{B}(Y)\cong M_{\rm Dol}(X{\setminus}0)$.